Which of the following statements is/are correct?

1.   The bond energy, BE(Cl-Cl), in Cl2 is 244 kJ mol-1 since ∆H‎°f {Br,atomic,g) = 122 kJ mol-1.

2.   A compound has a ∆H‎°f , calculated from average bond energies, of -20 kJ mol-1, whereas its experimental ∆H‎°f is -80 kJ mol-1. This compound is less stable than predicted by its Lewis structure.

3.   A gas phase chemical reaction is found to be exothermic. This means that the sum of the bond energies of the reactants is greater than the sum of the bond energies of the products.

  • A.

    1 only

  • B.

    none

  • C.

    3 only

  • D.

    2 only

  • E. 1 & 3 only
Answer #1

Answer is option A

1 is correct since dH ( Br atomic ) = 1/2 times of BE ( Cl-Cl) = 1/2( 244) = 122 KJmol-1

2 is false since experimenatl dH value showed - 80 KJ/mol. The lower the energy the more stable compound is. Hence compound is expermentally more stable than predicted

( as exprimental value of - 80 lower than predicted -20 )

3) Exothermic means heat is released in reaction or heat of reaction is -ve value

Heat of reaction = Bond energies of reactants - Bond energies of products

= +ve value ( since it was mentioned reactants bond energies higher than products so difference is +ve)

Thus based on bond energies we see we had endothermic ( heat absorbed ) but experments suggest exothermic. Hence bond energies of reactants are not greater than products.